OpenStudy (anonymous):

y varies indirectly as twice x. When x = 4, y = 1. Find y when x = 5. I think y = 2 is that correct?

OpenStudy (anonymous):

y = k/2x x=4,y=1 1 = k/8 k = 8 so y = 8/2x when x=8 y = 1/2